OpenStudy (anonymous):

how do you solve |3x+7| is greater than or equal to 26

OpenStudy (anonymous):

you are taking the absolute value of 3x+7| >= 26 so you write out the inequality and solve

OpenStudy (anonymous):

When you see I I, it means absolute value, or the distance from 0. This means the number isn't changing, but if it's negative, it will be a positive number instead. Absolute value always means a positive number. So, I3x+7I is greater or equal to (greater than sign with line under it) to 26. Whatever x is, you'd plug it in.

OpenStudy (kirbykirby):

\[|x|=\begin{cases} x &,\text{if }x\ge 0\\ -x&,\text{if }x<0\end{cases} \] Then: \[|3x+7|=\begin{cases} 3x+7 &,\text{if }3x+7\ge 0\\ -(3x+7)&,\text{if }3x+7<0\end{cases} \] So you have two cases to solve for: \((3x+7) \ge 26\) and \(-(3x+7)\ge 26\)